Two guns ,situated on the top of a hill of height 10 m ,fire one shot each with the same speed 5root 3 m/s at some interval of time .One gun fires horizontaly and other fires upwards at an angle of 60 with the horizontal .The shots collides in mid air at pt. P.            Find 1)the time interval btw firing  2)the coordinates of pt. P.(take origin at foot of hill  right below muzzles and trajectories in x-y plane).  pls give full sol.

Grade:12

2 Answers

AJIT AskiitiansExpert-IITD
13 years ago

Dear jyoti ,

write equations of motion for both the guns

gun2:fires at t=0

x direction : x =5root3 * cos 60 * (t )   

y direction : y =  5root3sin60(t) - 1/2 *g *(t)2

gun1:fires at t=t1

x direction : x =5root3 *(t-t1)   

y direction : y =  - 1/2 *g *(t-t1)2   (negative as upward direction is positive


 

when they collide their x and y cordinates are equal ,

so if  they collide at t=t2 then ,

5root3 *t2-t1= 5root3 * cos 60 * (t2 )   

- 1/2 *g *(t2-t1)2  =  5root3sin60(t2) - 1/2 *g *(t2)2  

solving we get ,

t2= 2s , t1 =1s

so 1) time interval b/w firing  =t1 =1s

   2)p = (5root3 *(2-1) ,  -1/2*g*(2-1)2 )

      p =(5root3 , -5)

                                                                    
                                                                     
Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.


Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Win exciting gifts by answering the questions on Discussion Forum..

Hrishi Raaj Singh Chauhan
6 years ago
Find trajectory eqn of both then equate it then solve it. Then u will get x and then u can find t1 and t2